The "human cannonball" has long been a popular and extremely dangerous circus

stunt.

A person with a mass of 45kg starts at rest and leaves the cannon with a final speed of

25m/s after a 1600N force is applied to them by the cannon. For how much time was the

force being applied to the person?


1
Expert's answer
2022-04-11T12:07:54-0400

The time of force acting

"\Delta t=\frac{m\Delta v}{F}"

"\Delta t=\frac{45*25}{1600}=0.703\:\rm s"